OpenStudy (itsbribro):

How many solutions does the equation have? 3(4x-2)=2(6x-3) A. one solution B. infinite solutions C. no solution

OpenStudy (gorv):

12x-6=12x-6 both side are same

OpenStudy (gorv):

so any value of x satisfy this means infinite solution

OpenStudy (gorv):

you can put any value of x

OpenStudy (gorv):

throw it here :P

OpenStudy (itsbribro):

How many solutions does the equation have? \[6-a=\frac{ 1 }{ 3 }(9a-18)\] A. one solution B. infinite solutions C. no solution

OpenStudy (gorv):

multiply both side by 3 6-a=(3*(3a-6))/3=3a-6 6-a=3a-6 6+6=3a+a 12=4a a=3